4. A customer wants to make a teepee in his backyard for his children. He plans to use lengths of PVC plumbing pipe for the supports on the teepee, and he wants the teepee to be 12 feet across and 8 feet tall . How long should the PVC plumbing pipe be?


Answer by checkley71(8403)

12 FEET IS THE BASE. SO WE CONSTRUCT A RIGHT TRIANGLE WITH A BASE=12/2=6 FEET & ANOTHER SIDE 8 FEET. NOW WE NEED TO FIND THE HYPOTENUSE (C).
6^2+8^2=C^2
36=64=C^2
100=C^2
C=SQRT100
C=10 FEET FOR THE POLE LENGTH WILL DO THE JOB.